### v2

# The Cayley-Bacharach property and the Levinson-Ullery conjecture

Abstract.
In this paper, we study the geometric configurations of a finite set of points having the Cayley-Bacharach property in the $n$-dimensional projective space ${\mathbb{P}}^{n}$. Our main contribution is the proof of the Levinson-Ullery conjecture for the previously unsolved case where $d=4$ and $r\geq 1$.
Key words and phrases:
Projective geometry, Cayley–Bacharach property, plane arrangement
2020 Mathematics Subject Classification:
Primary 14N05; Secondary 14N20, 14N25
1. Introduction
Let $\mathbb{X}$} be a finite set of points in the $n$-dimensional projective space ${\mathbb{P}}^{n}$ over a field of characteristic zero. We are interested in studying the intrinsic geometric properties of $\mathbb{X}$}, particularly, the Cayley-Bacharach property. This property is a geometric invariant that has attracted considerable and ongoing research interest. Originating in classical projective geometry, the Cayley-Bacharach property has a long and rich history and remains a key concept in modern algebraic geometry. Geometrically, the finite set of points $\mathbb{X}$} is said to have the Cayley-Bacharach property of degree $r$ (for short, $\mathbb{X}$} has $\mathop{\rm CBP}\nolimits(r)$) if every hypersurface of degree $r$ that contains all but one point of $\mathbb{X}$} automatically contains the last point. The most celebrated classical result concerning this notion is the Cayley-Bacharach theorem, which asserts that if $\mathbb{X}$} is the complete intersection of two plane curves of degrees $d$ and $e$, respectively, then $\mathbb{X}$} has $\mathop{\rm CBP}\nolimits(d+e-3)$ (see [2, 3]). Since then, this result has been generalized extensively to various contexts (see, e.g., [4, 5, 6, 9, 10, 11, 13, 14]). Furthermore, the property has not only led to the formulation of many interesting conjectures (see [5, 16] for historical background and open problems), but has also yielded significant applications to different branches of mathematics. These applications include the algebraic characterizations of specific 0-dimensional schemes like complete intersections and arithmetically Gorenstein schemes [12, 14, 15], its use in studying the measures of irrationality for projective varieties [1, 17, 18], and its crucial role in coding theory [7, 8].
In this paper we focus on investigating the geometrical configurations of a finite set of points ${\mathbb{X}}\subseteq{\mathbb{P}}^{n}$ satisfying the Cayley-Bacharach property. An interesting result on this direction is the following theorem (see [1, Lemma 2.4]).
Theorem 1.1.
Let ${\mathbb{X}}\subseteq{\mathbb{P}}^{n}$ be a finite set of points having $\mathop{\rm CBP}\nolimits(r)$. If
$|{\mathbb{X}}|\leq 2r+1$
then $\mathbb{X}$} lies on a line.
Inspired by this result, in [16], Levinson and Ullery introduced the notion of a plane configuration to describe the geometry of $\mathbb{X}$} and proposed the following conjecture. Here a plane configuration $\mathcal{P}$ is a union of distinct positive-dimensional linear spaces $\mathcal{P}=\bigcup_{i=1}^{k}\mathcal{P}_{i}$ in ${\mathbb{P}}^{n}$, its dimension is $\dim(\mathcal{P})=\sum_{i=1}^{k}\dim(\mathcal{P}_{i})$ and its length is $l(\mathcal{P})=k$.
Conjecture 1.2 (Levinson-Ullery).
Let ${\mathbb{X}}\subseteq{\mathbb{P}}^{n}$ be a set of points having $\mathop{\rm CBP}\nolimits(r)$. If
$|{\mathbb{X}}|\leq(d+1)r+1$
then $\mathbb{X}$} lies on a plane configuration $\mathcal{P}$ of dimension $d.$
Clearly, Conjecture 1.2 holds true for $d=1$ by Theorem 1.1. The main result established in the work of Levinson and Ullery [16] is summarized as follows (see [16, Theorem 1.3]):
Theorem 1.3.
The conjecture 1.2 holds in the following cases:
  1. (1)
    For all pairs $(d,r)$ with $r\leq 2$ and $d\geq 1$.
  2. (2)
    For all pairs $(d,r)$ with $d\leq 3$ and $r\geq 1$.
  3. (3)
    For the pair $(d,r)=(4,3)$.
It is therefore natural to investigate whether Conjecture 1.2 remains true for all pairs $(d,r)$ with $d\geq 4$ and $r\geq 3$. We address the case where $d=4$ and $r\geq 1$ and provide the following main result.
Theorem 1.4.
The Conjecture 1.2 holds for all pairs $(d,r)$ with $d=4$ and $r\geq 1$.
The paper is structured as follows. In Section 2, we establish the necessary background. We recall the definitions of the Cayley-Bacharach property, Hilbert functions, canonical modules, and plane configurations. We then present several algebraic characterizations of the Cayley-Bacharach property. Additionally, we include technical results based on the work of [16]. Section 3 is devoted to the proof of the main theorem and a requisite proposition.
2. Preliminary
In what follows we work over a field $K$ of characteristic zero. By ${\mathbb{P}}^{n}$ we denote the $n$-dimensional projective space over $K$. Our main objects of study are finite sets of points ${\mathbb{X}}\subseteq{\mathbb{P}}^{n}$. Let $I_{\mathbb{X}}$ denote the homogeneous vanishing ideal of $\mathbb{X}$} in $P=K[X_{0},\dots,X_{n}]$, and let $R=P/I_{\mathbb{X}}$ be the homogeneous coordinate ring of $\mathbb{X}$}. The ring $R$ is a 1-dimensional standard graded $K$-algebra. After performing a homogeneous linear change of coordinates (see, e.g., [10, Lemma 6.3.20]), we may assume that $x_{0}$, the image of $X_{0}$ in $R$, is a non-zerodivisor of $R$.
The Hilbert function of $\mathbb{X}$} is the function $\mathop{\rm HF}\nolimits_{\mathbb{X}}:{\mathbb{Z}}\rightarrow{\mathbb{N}},$ $i\mapsto\dim_{K}(R_{i})$, and its first difference function $\Delta\mathop{\rm HF}\nolimits_{\mathbb{X}}(i):=\mathop{\rm HF}\nolimits_{\mathbb{X}}(i)-\mathop{\rm HF}\nolimits_{\mathbb{X}}(i-1)$. We have $\mathop{\rm HF}\nolimits_{\mathbb{X}}(i)=0$ for $i<0$ and
$1=\mathop{\rm HF}\nolimits_{\mathbb{X}}(0)<\mathop{\rm HF}\nolimits_{\mathbb{X}}(1)<\cdots<\mathop{\rm HF}\nolimits_{\mathbb{X}}(r_{\mathbb{X}})=\mathop{\rm HF}\nolimits_{\mathbb{X}}(r_{\mathbb{X}}+1)=\cdots=|{\mathbb{X}}|,$
where $r_{\mathbb{X}}$ is the regularity index of $\mathbb{X}$}.
Definition 2.1.
For $r\geq 0$, we say that $\mathbb{X}$} has the Cayley-Bacharach property of degree $r$ (for short, $\mathbb{X}$} has $\mathop{\rm CBP}\nolimits(r)$) if every hypersurface of degree $r$ which contains all but one point of $\mathbb{X}$} must contain all the points of $\mathbb{X}$}. In the case that $\mathbb{X}$} has $\mathop{\rm CBP}\nolimits(r_{\mathbb{X}}-1)$ we also say that $\mathbb{X}$} is a Cayley-Bacharach scheme.
Notice that $\mathbb{X}$} has $\mathop{\rm CBP}\nolimits(r)$ if and only if for any $p\in{\mathbb{X}}$ we have $\mathop{\rm HF}\nolimits_{{\mathbb{X}}\setminus\{p\}}(r)=\mathop{\rm HF}\nolimits_{\mathbb{X}}(r)$. If $\mathbb{X}$} has $\mathop{\rm CBP}\nolimits(r)$ then it has $\mathop{\rm CBP}\nolimits(r-1)$. Moreover, the number $r_{\mathbb{X}}-1$ is the largest degree $r\geq 0$ such that $\mathbb{X}$} can have $\mathop{\rm CBP}\nolimits(r)$.
Remark 2.2.
Let $\mathbb{Y}$} be a subset of $\mathbb{X}$} with $|{\mathbb{Y}}|=|{\mathbb{X}}|-1$.
  1. (a)
    The image of the vanishing ideal $I_{\mathbb{Y}}\subseteq P$ of $\mathbb{Y}$} in $R=P/I_{\mathbb{X}}$ is denoted by $I_{{\mathbb{Y}}/{\mathbb{X}}}$. Its initial degree is $\alpha_{{\mathbb{Y}}/{\mathbb{X}}}:=\min\{k\in{\mathbb{N}}\mid(I_{{\mathbb{Y}}/{\mathbb{X}}})_{k}\neq 0\}$. Then $\alpha_{{\mathbb{Y}}/{\mathbb{X}}}\leq r_{\mathbb{X}}$ and the Hilbert function of $\mathbb{Y}$} satisfies
    $\mathop{\rm HF}\nolimits_{{\mathbb{Y}}}(i)\;=\;\begin{cases}\mathop{\rm HF}\nolimits_{\mathbb{X}}(i)&\text{for $i<\alpha_{{\mathbb{Y}}/{\mathbb{X}}}$},\\ \mathop{\rm HF}\nolimits_{\mathbb{X}}(i)-1&\text{for $i\geq\alpha_{{\mathbb{Y}}/{\mathbb{X}}}$}.\end{cases}$
  2. (b)
    A nonzero homogeneous element $f^{*}_{\mathbb{Y}}\in(I_{{\mathbb{Y}}/{\mathbb{X}}})_{\alpha_{{\mathbb{Y}}/{\mathbb{X}}}}$ is called a minimal separator of $\mathbb{Y}$} in $\mathbb{X}$}. In this case, $(I_{{\mathbb{Y}}/{\mathbb{X}}})_{\alpha_{{\mathbb{Y}}/{\mathbb{X}}}+i}=Kx_{0}^{i}f^{*}_{\mathbb{Y}}$ for every $i\geq 0$.
  3. (c)
    The graded $R$-module $\omega_{R}=\mathop{\rm Hom}\nolimits_{K[x_{0}]}(R,K[x_{0}])(-1)$ is called the canonical module of $R$. Its $R$-module structure is defined by $(f\cdot\varphi)(g)=\varphi(fg)$ for all $f,g\in R$ and $\varphi\in\omega_{R}$. It is also a finitely generated graded $R$-module and for every $i\in{\mathbb{Z}}$ we have
    $\mathop{\rm HF}\nolimits_{\omega_{R}}(i)=|{\mathbb{X}}|-\mathop{\rm HF}\nolimits_{\mathbb{X}}(-i)$
    (see [9, Lemma 1.3]).
From [6, Proposition 2.1] and [13, Proposition 4.3] we have the following characterizations of the Cayley-Bacharach property of $\mathbb{X}$}.
Proposition 2.3.
The following conditions are equivalent for every $r\geq 1$:
  1. (a)
    $\mathbb{X}$} has $\mathop{\rm CBP}\nolimits(r)$.
  2. (b)
    Every subset $\mathbb{Y}$} of $\mathbb{X}$} with $|{\mathbb{Y}}|=|{\mathbb{X}}|-1$ satisfies $\alpha_{{\mathbb{Y}}/{\mathbb{X}}}\geq r+1$.
  3. (c)
    No element of $(I_{{\mathbb{Y}}/{\mathbb{X}}})_{r_{\mathbb{X}}}\setminus\{0\}$ is divisible by $x_{0}^{r_{\mathbb{X}}-r}$
  4. (d)
    There exists an element $\varphi\in(\omega_{R})_{-r}$ such that $\mathop{\rm Ann}\nolimits_{R}(\varphi)=0$.
Corollary 2.4.
Let ${\mathbb{X}}\subseteq{\mathbb{P}}^{n}$ be a finite set of points having $\mathop{\rm CBP}\nolimits(r)$.
  1. (a)
    $\mathop{\rm HF}\nolimits_{\mathbb{X}}(i)+\mathop{\rm HF}\nolimits_{\mathbb{X}}(r-i)\leq|X|$ for all $i\in{\mathbb{Z}}$.
  2. (b)
    We have $|{\mathbb{X}}|\geq r+2$.
Proof.
By Proposition 2.3(d), there exists $\varphi\in(\omega_{R})_{-r}$ such that $\mathop{\rm Ann}\nolimits_{R}(\varphi)=0$. The multiplication map $R\rightarrow R\cdot\varphi\subseteq\omega_{R}(-r),f\mapsto f\cdot\varphi$ is injective, and hence
$\mathop{\rm HF}\nolimits_{\mathbb{X}}(i)\leq\mathop{\rm HF}\nolimits_{\omega_{R}(-r)}(i)=\mathop{\rm HF}\nolimits_{\omega_{R}}(i-r)=|{\mathbb{X}}|-\mathop{\rm HF}\nolimits_{\mathbb{X}}(r-i)$
for all $i\in{\mathbb{Z}}$. This completes the proof of (a). Claim (b) follows from (a) or [1, Lemma 2.4]. ∎
Now we recall from [16, Definition 2.1] the following notion of plane configurations.
Definition 2.5.
A plane configuration $\mathcal{P}$ is a union of distinct positive-dimensional linear spaces $\mathcal{P}=\bigcup_{i=1}^{k}\mathcal{P}_{i}$ in ${\mathbb{P}}^{n}$. The number $\dim(\mathcal{P}):=\sum_{i=1}^{k}\dim(\mathcal{P}_{i})$ is called the dimension of $\mathcal{P}$ and $\ell(\mathcal{P}):=k$ is called the length of $\mathcal{P}$.
Note that a 0-dimensional plane configuration is empty, a 1-dimensional plane configuration is a single line, and a 2-dimensional plane configuration is either two lines or one 2-plane. Here a $d$-plane refers to a $d$-dimensional linear space. Also, the linear span of a plane configuration $\mathcal{P}=\bigcup_{i=1}^{k}\mathcal{P}_{i}$ is denoted by $\mathop{\rm span}\nolimits(\mathcal{P})$ or $\mathop{\rm span}\nolimits(\mathcal{P}_{i}\mid i=1,\dots,k)$.
Definition 2.6.
Let $\mathcal{P}=\bigcup_{i=1}^{k}\mathcal{P}_{i}$ be a plane configuration in ${\mathbb{P}}^{n}$.
  1. (a)
    $\mathcal{P}$ is said to be skew if $\mathcal{P}_{i}\cap\mathcal{P}_{j}=\emptyset$ for all $i\neq j$.
  2. (b)
    $\mathcal{P}$ is said to be split if $\mathcal{P}_{i}\cap\mathop{\rm span}\nolimits(\mathcal{P}_{j}\mid j\neq i)=\emptyset$ for every $i=1,\dots,k$.
Clearly, if $\mathcal{P}$ is split then it is skew, and the converse holds true if $\ell(\mathcal{P})=2$. Also, $\mathcal{P}$ is split if and only if $\dim(\mathop{\rm span}\nolimits(\mathcal{P}))=\dim(\mathcal{P})+\ell(\mathcal{P})-1$.
The following proposition follows from [16, Proposition 2.5].
Proposition 2.7.
Let ${\mathbb{X}}\subseteq{\mathbb{P}}^{n}$ be a finite set of points, and let $\mathcal{P}$ be a plane configuration of length $\ell(\mathcal{P})=k$. If $\mathbb{X}$} has $\mathop{\rm CBP}\nolimits(r)$ then ${\mathbb{X}}\setminus\mathcal{P}$ has $\mathop{\rm CBP}\nolimits(r-k)$.
The results presented in [16, Propositions 4.2, 4.7 and Lemma 4.5(1)] lead to the next proposition.
Proposition 2.8.
Let ${\mathbb{X}}\subseteq{\mathbb{P}}^{n}$ be a finite set of points, and let $\mathcal{P}=\bigcup_{i=1}^{k}\mathcal{P}_{i}$ be a plane configuration of length $\ell(\mathcal{P})=k$ such that ${\mathbb{X}}_{\mathcal{P}_{i}}:={\mathbb{X}}\cap\mathcal{P}_{i}\neq\emptyset$ for all $i=1,\dots,k$.
  1. (a)
    If $\mathcal{P}$ is split and ${\mathbb{X}}\subseteq\mathcal{P}$, then $\mathbb{X}$} has $\mathop{\rm CBP}\nolimits(r)$ if and only if, for each $i$, ${\mathbb{X}}\cap\mathcal{P}_{i}$ has $\mathop{\rm CBP}\nolimits(r)$.
  2. (b)
    If $\mathbb{X}$} has $\mathop{\rm CBP}\nolimits(r)$ and $\mathcal{P}$ is skew, then one of the following holds:
    1. (i)
      Each plane $\mathcal{P}_{i}$ contains at least $\max\{k,r+2)$ points of $\mathbb{X}$}, or
    2. (ii)
      Some plane $\mathcal{P}_{i}$ contains fewer than $k$ points, and also $k\geq r+2.$
  3. (c)
    Suppose that $\mathbb{X}$} has $\mathop{\rm CBP}\nolimits(r)$, $k=2$, and $\mathcal{P}_{1}$ and $\mathcal{P}_{2}$ meet at a point $p\in{\mathbb{P}}^{n}$. Then at least one of ${\mathbb{X}}_{\mathcal{P}_{1}}$ and ${\mathbb{X}}_{\mathcal{P}_{1}}\cup\{p\}$ has $\mathop{\rm CBP}\nolimits(r).$
3. Proof of main theorem
As in the previous section, let $r\geq 1$ and let ${\mathbb{X}}\subseteq{\mathbb{P}}^{n}$ be a finite set of points having $\mathop{\rm CBP}\nolimits(r)$. This section is devoted to the proof of Theorem 1.4, which states that if $|{\mathbb{X}}|\leq 5r+1$ then $\mathbb{X}$} lies on a plane configuration $\mathcal{P}$ of dimension $4$. This proof will utilize the following proposition.
Proposition 3.1.
Let ${\mathbb{X}}\subseteq{\mathbb{P}}^{n}$ be a finite set of points having $\mathop{\rm CBP}\nolimits(r)$ and $|{\mathbb{X}}|\leq(d+1)r+1$, let $A$ be a $k$-plane with $k\leq d$, and let ${\mathbb{X}}_{A}:={\mathbb{X}}\cap A$ and ${\mathbb{X}}_{B}:={\mathbb{X}}\setminus{\mathbb{X}}_{A}$.
  1. (a)
    If Conjecture 1.2 holds for (d1,r)(d-1,r) and $\mathbb{X}$} does not lie on a plane configuration of dimension d1d-1, then $|{\mathbb{X}}|\geq dr+2$.
  2. (b)
    Assume that Conjecture 1.2 holds for every pair $(i,j)$ with $i\leq d$ and $j<r$.
    1. (i)
      If $|{\mathbb{X}}_{A}|\geq d+1$ then ${\mathbb{X}}_{B}$ lies on a plane configuration of dimension $d$.
    2. (ii)
      If ${\mathbb{X}}_{A}$ does not lie on a plane configuration of dimension (d1)(d-1) and ${\mathbb{X}}_{B}$ is contained in an $\ell$-plane with $\ell\leq d$ and $d\leq r$, then ${\mathbb{X}}={\mathbb{X}}_{A}$.
Proof.
(a) Note that $\mathbb{X}$} has $\mathop{\rm CBP}\nolimits(r)$. If $|{\mathbb{X}}|\leq dr+1$, then the validity of Conjecture 1.2 for the pair (d1,r)(d-1,r) guarantees that $\mathbb{X}$} lies on a plane configuration of dimension d1d-1.
(b) If ${\mathbb{X}}_{B}=\emptyset$ then we are done. Suppose that ${\mathbb{X}}_{B}\neq\emptyset$. By Proposition 2.7, ${\mathbb{X}}_{B}$ has $\mathop{\rm CBP}\nolimits(r-1)$. Since $|{\mathbb{X}}_{A}|\geq d+1$, we have
$1\leq|{\mathbb{X}}_{B}|\leq(d+1)r+1-(d+1)=(d+1)(r-1)+1.$
Because Conjecture (b) holds true for (d,r1)(d,r-1), the set ${\mathbb{X}}_{B}$ must lie on a plane configuration of dimension $d$. Thus claim (i) follows.
Next, we prove (ii). Observe that $A$ is a $d$-plane and $|{\mathbb{X}}_{A}|\geq d+1$. Suppose for a contradiction that ${\mathbb{X}}_{B}\neq\emptyset$. Let $B$ be an $\ell$-plane containing ${\mathbb{X}}_{B}$. By Proposition 2.7, both ${\mathbb{X}}_{A}$ and ${\mathbb{X}}_{B}$ have $\mathop{\rm CBP}\nolimits(r-1)$. Since ${\mathbb{X}}_{A}$ does not lie on a plane configuration of dimension (d1)(d-1) and Conjecture 1.2 holds for $(i,j)$ with $i\leq d$ and $j\leq r-1$, claim (a) yields
$|{\mathbb{X}}_{A}|\;\geq\;d(r-1)+2.$
It follows that
$|{\mathbb{X}}_{B}|\leq((d+1)r+1)-(d(r-1)+2)=r+d-1\leq 2(r-1)+1,$
and consequently ${\mathbb{X}}_{B}$ lies on a line $L$ by Theorem 1.1. Now we distinguish the following two cases.
  1. (i)
    AA and $L$ are split. In this case, ${\mathbb{X}}_{A}$ and ${\mathbb{X}}_{B}$ have $\mathop{\rm CBP}\nolimits(r)$ by Proposition 2.8(a). Due to Corollary 2.4 and (a), the lower bounds for the cardinalities of ${\mathbb{X}}_{A}$ and ${\mathbb{X}}_{B}$ are given by
    $|{\mathbb{X}}_{A}|\geq dr+2\quad\text{and}\quad|{\mathbb{X}}_{B}|\geq r+2.$
    This implies $(d+1)r+4\leq|{\mathbb{X}}_{A}|+|{\mathbb{X}}_{B}|=|{\mathbb{X}}|\leq(d+1)r+1$, which is impossible.
  2. (ii)
    AA and $L$ intersect at a point, say $p$. By Lemma 2.8(c), at least one of ${\mathbb{X}}_{A}$ and ${\mathbb{X}}_{A}\cup\{p\}$ has $\mathop{\rm CBP}\nolimits(r).$ If ${\mathbb{X}}_{A}$ has $\mathop{\rm CBP}\nolimits(r)$, then claim (a) yields that $|{\mathbb{X}}_{A}|\geq dr+2$. When ${\mathbb{X}}_{A}\cup\{p\}$ has $\mathop{\rm CBP}\nolimits(r),$ then $|{\mathbb{X}}_{A}\cup\{p\}|\geq dr+2$ by (a). In these both cases, we always have $|{\mathbb{X}}_{A}|\geq dr+1$. On the other hand, the set ${\mathbb{X}}_{B}$ has $\mathop{\rm CBP}\nolimits(r-1)$, and Corollary 2.4 shows that $|{\mathbb{X}}_{B}|\geq r+1$. Hence, we obtain $|{\mathbb{X}}|=|{\mathbb{X}}_{A}|+|{\mathbb{X}}_{B}|\geq(dr+1)+(r+1)\geq(d+1)r+2$, a contradiction.
Altogether, we find ${\mathbb{X}}_{B}=\emptyset$ and ${\mathbb{X}}={\mathbb{X}}_{A}$. ∎
Now we are ready to give a proof of our main result.
Proof of Theorem 1.4.
According to Theorem 1.3, it suffices to consider the case $r\geq 4$ and use induction on $r$. Let ${\mathbb{X}}\subseteq{\mathbb{P}}^{n}$ be a finite set of points having $\mathop{\rm CBP}\nolimits(r)$ and $|{\mathbb{X}}|\leq 5r+1$. Further, let $A$ be a 4-plane containing the maximum number of points of $\mathbb{X}$}, let ${\mathbb{X}}_{A}:={\mathbb{X}}\cap A$, and let ${\mathbb{X}}_{B}:={\mathbb{X}}\setminus{\mathbb{X}}_{A}$. When ${\mathbb{X}}_{B}=\emptyset$, the theorem holds true. So, we consider the case ${\mathbb{X}}_{B}\neq\emptyset$. In this case $|{\mathbb{X}}_{A}|\geq 5$, and Proposition 3.1(b.i) yields that the set ${\mathbb{X}}_{B}$ lies on a plane configuration of dimension 4. Also, the set ${\mathbb{X}}_{B}$ has $\mathop{\rm CBP}\nolimits(r-1)$ by Proposition 2.7. In the following we consider different cases depending on the geometric configurations of ${\mathbb{X}}_{B}$.
  1. Case 1:
    ${\mathbb{X}}_{B}$ lies on a 4-plane.
    If ${\mathbb{X}}_{A}$ does not lie on a plane configuration of dimension 3, then, by the inductive hypothesis, Proposition 3.1(b.ii) shows that ${\mathbb{X}}={\mathbb{X}}_{A}$, and subsequently $\mathbb{X}$} lies on a plane configuration of dimension 4, as wanted. By the choice of $A$, ${\mathbb{X}}_{A}$ lies on a 4-plane, but not a 3-plane. Thus, it suffices to look at the case that ${\mathbb{X}}_{A}$ lies on a plane configuration of dimension 3, i.e., it lies on either a skew plane configuration of three lines or a split plane configuration of a 2-plane and a line.
    1. (1.1)
      ${\mathbb{X}}_{A}$ lies on a skew plane configuration of three lines $L_{1},L_{2},L_{3}$.
      Note that ${\mathbb{X}}_{A}$ has $\mathop{\rm CBP}\nolimits(r-1)$ by Proposition 2.7 and $r\geq 4$. By Proposition 2.8(b), each line $L_{i}$ contains at least $r+1$ points of ${\mathbb{X}}_{A}$ for $i=1,2,3$. Consequently, we have $|{\mathbb{X}}_{A}|\geq 3(r+1)$. It follows that
      $|{\mathbb{X}}_{B}|\leq 5r+1-3(r+1)=2r-2<2(r-1)+1.$
      According to Theorem 1.1, the set ${\mathbb{X}}_{B}$ lies on a line, and therefore, the set $\mathbb{X}$} lies on a plane configuration of dimension 4, as desired.
    2. (1.2)
      ${\mathbb{X}}_{A}$ lies on a split plane configuration of a 2-plane $H$ and a line $L$.
      Since ${\mathbb{X}}_{A}$ has $\mathop{\rm CBP}\nolimits(r-1)$, Proposition 2.8(a) implies that two sets of points ${\mathbb{X}}_{A}\cap H$ and ${\mathbb{X}}_{A}\cap L$ also have $\mathop{\rm CBP}\nolimits(r-1)$. Moreover, ${\mathbb{X}}_{A}\cap H$ does not lie on a line by the choice of $A$. By Corollary 2.4 and Proposition 3.1(a), we get
      $|{\mathbb{X}}_{A}\cap L|\geq r+1\quad\text{and}\quad|{\mathbb{X}}_{A}\cap H|\geq 2(r-1)+2=2r.$
      This implies
      $|{\mathbb{X}}_{B}|\leq 5r+1-(r+1)-2r\leq 2(r-1)+2.$
      1. (i)
        If $|{\mathbb{X}}_{B}|\leq 2(r-1)+1$ then ${\mathbb{X}}_{B}$ must lies on a line by Theorem 1.1, and hence the set $\mathbb{X}$} is contained in a plane configuration of dimension 4, as wanted to show.
      2. (ii)
        If $|{\mathbb{X}}_{B}|=2r\leq 3(r-1)+1$, then ${\mathbb{X}}_{B}$ lies on a plane configuration of dimension 2 by the inductive hypothesis and $|{\mathbb{X}}_{A}|=3r+1$. When ${\mathbb{X}}_{B}$ lies on a line, as above, the set $\mathbb{X}$} lies on a plane configuration of dimension 4 and we are done. If ${\mathbb{X}}_{B}$ lies on a union of two split lines, then each of these lines contains at least $r+1$ points of ${\mathbb{X}}_{B}$ by Proposition 2.8(a) and Corollary 2.4, and so $|{\mathbb{X}}_{B}|\geq 2r+2$, which is impossible. Here we remark that a skew plane configuration of two lines agrees with a split plane configuration of two lines. Now we consider the last subcase that the set ${\mathbb{X}}_{B}$ lies on a $2$-plane $H^{\prime}$. If $H^{\prime}\cap H\neq\emptyset$ or $H^{\prime}\cap L\neq\emptyset$ then the pair $(H,H^{\prime})$ defines a 4-plane containing at least $4r$ points of $\mathbb{X}$} or the pair $(L,H^{\prime})$ and a point in ${\mathbb{X}}_{A}\cap H$ define a 4-plane containing at least $3r+2$ points of $\mathbb{X}$}, which contradicts to the choice of the 4-plane $A$. Hence we must have $H^{\prime}\cap H=H^{\prime}\cap L=\emptyset$. In particular, $\mathbb{X}$} is contained in a skew plane configuration $\mathcal{P}=H\cup L\cup H^{\prime}$ of length $\ell(\mathcal{P})=3<r$. Since $\mathbb{X}$} has $\mathop{\rm CBP}\nolimits(r)$, an application of Proposition 2.8(b) yields that the line $L$ contains at least $r+2$ points of $\mathbb{X}$}. It follows that
        $|{\mathbb{X}}|=|{\mathbb{X}}\cap H|+|{\mathbb{X}}\cap L|+|{\mathbb{X}}\cap H^{\prime}|\geq 2r+(r+2)+2r=5r+2.$
        However, this contradicts the assumption that $|{\mathbb{X}}|\leq 5r+1$.
Next, we assume that no 4-plane contains ${\mathbb{X}}_{B}$ and we consider the following remaining cases.
  1. Case 2:
    ${\mathbb{X}}_{B}$ is contained in the union of a 3-plane $H$ and a line $L$.
    Since ${\mathbb{X}}_{B}$ is not contained in a $4$-plane, then the 3-plane $H$ and the line $L$ must be split. We define two subsets of ${\mathbb{X}}_{B}$ by ${\mathbb{X}}_{H}:={\mathbb{X}}_{B}\cap H$ and ${\mathbb{X}}_{L}:={\mathbb{X}}_{B}\cap L$. Remark that ${\mathbb{X}}_{B}$ has $\mathop{\rm CBP}\nolimits(r-1)$. By Proposition 2.8(a), two subsets ${\mathbb{X}}_{H}$ and ${\mathbb{X}}_{L}$ also have $\mathop{\rm CBP}\nolimits(r-1)$. By Corollary 2.4, we have $|{\mathbb{X}}_{L}|\geq r+1$. Observe that ${\mathbb{X}}_{H}$ may lie on a plane configuration of dimension 2, but it does not lie on a line. By Proposition 3.1(a), the cardinality of ${\mathbb{X}}_{H}$ is bounded below as
    $|{\mathbb{X}}_{H}|\geq 2(r-1)+2=2r.$
    Moreover, since the 4-plane $A$ contains the maximum number of points in $\mathbb{X}$}, we must have
    $|{\mathbb{X}}_{A}|\geq|{\mathbb{X}}_{H}|+1\geq 2r+1.$
    Thus, combining the lower bounds for the disjoint sets ${\mathbb{X}}_{A},{\mathbb{X}}_{H},{\mathbb{X}}_{L}$, we find
    $|{\mathbb{X}}|=|{\mathbb{X}}_{A}|+|{\mathbb{X}}_{H}|+|{\mathbb{X}}_{L}|\geq(2r+1)+2r+r+1=5r+2,$
    which contradicts the hypothesis that $|{\mathbb{X}}|\leq 5r+1$.
  2. Case 3:
    ${\mathbb{X}}_{B}$ is contained in the union of two 2-planes $K_{1}$ and $K_{2}$ such that ${\mathbb{X}}_{K_{i}}:={\mathbb{X}}_{B}\cap K_{i}$ does not lie on a line for $i=1,2$.
    Clearly, the two planes $K_{1}$ and $K_{2}$ are split, as ${\mathbb{X}}_{B}$ is not contained in a $4$-plane. Also, we see that ${\mathbb{X}}_{B}={\mathbb{X}}_{K_{1}}\cup{\mathbb{X}}_{K_{2}}$ and ${\mathbb{X}}_{B}$ has $\mathop{\rm CBP}\nolimits(r-1)$. It follows from Proposition 2.8(a) that both ${\mathbb{X}}_{K_{1}}$ and ${\mathbb{X}}_{K_{2}}$ have $\mathop{\rm CBP}\nolimits(r-1)$. Thank to Proposition 3.1(a), the cardinality of ${\mathbb{X}}_{B}$ satisfies
    $|{\mathbb{X}}_{B}|=|{\mathbb{X}}_{K_{1}}|+|{\mathbb{X}}_{K_{2}}|\geq 2r+2r=4r.$
    By the choice of the 4-plane $A$, we must have
    $|{\mathbb{X}}_{A}|\geq|{\mathbb{X}}_{K_{1}}|+2\geq 2r+2.$
    Hence we get $|{\mathbb{X}}|=|{\mathbb{X}}_{A}|+|{\mathbb{X}}_{B}|\geq 4r+2r+2=6r+2>5r+1$, which is impossible.
  3. Case 4:
    ${\mathbb{X}}_{B}$ is contained in the union of a $2$-plane $K$ and two lines $L_{1}$ and $L_{2}$ such that ${\mathbb{X}}_{K}:={\mathbb{X}}_{B}\cap K$ does not lie on a line.
    In this case $L_{1}$ and $L_{2}$ must be skew, since otherwise we would go back to Case 3 by defining the 2-plane $K^{\prime}=\mathop{\rm span}\nolimits(L_{1},L_{2})$. If $K\cap L_{i}\neq\emptyset$ for $i\in\{1,2\}$ then we go back to Case 2. Thus $K\cup L_{1}\cup L_{2}$ is a skew plane configuration containing ${\mathbb{X}}_{B}$ and ${\mathbb{X}}_{B}={\mathbb{X}}_{K}\cup{\mathbb{X}}_{L_{1}}\cup{\mathbb{X}}_{L_{2}}$, where ${\mathbb{X}}_{L_{1}}:={\mathbb{X}}_{B}\cap L_{1}$ and ${\mathbb{X}}_{L_{2}}:={\mathbb{X}}_{B}\cap L_{2}$. The linear space $H:=\mathop{\rm span}\nolimits(L_{1},L_{2})$ is a $3$-plane. Put ${\mathbb{X}}_{H}:={\mathbb{X}}_{B}\cap H$. We distinguish the following subcases.
    1. (4.1)
      $H\cap K=\emptyset$: By Proposition 2.8(a), the sets ${\mathbb{X}}_{H}$ and ${\mathbb{X}}_{K}$ have $\mathop{\rm CBP}\nolimits(r-1)$, since ${\mathbb{X}}_{B}$ has $\mathop{\rm CBP}\nolimits(r-1)$. From ${\mathbb{X}}_{H}={\mathbb{X}}_{L_{1}}\cup{\mathbb{X}}_{L_{2}}$ and $L_{1}\cap L_{2}=\emptyset$, again Proposition 2.8(a) yields that ${\mathbb{X}}_{L_{1}}$ and ${\mathbb{X}}_{L_{2}}$ have $\mathop{\rm CBP}\nolimits(r-1)$. By Corollary 2.4 and Proposition 3.1(a), we have
      $|{\mathbb{X}}_{K}|\geq 2(r-1)+2,\quad|{\mathbb{X}}_{L_{1}}|\geq r+1,\quad|{\mathbb{X}}_{L_{2}}|\geq r+1,$
      and consequently
      $|{\mathbb{X}}_{B}|=|{\mathbb{X}}_{K}|+|{\mathbb{X}}_{L_{1}}|+|{\mathbb{X}}_{L_{2}}|\geq 2r+2(r+1)=4r+2.$
      Furthermore, by the choice of the 4-plane $A$, we find
      $|{\mathbb{X}}_{A}|\geq|{\mathbb{X}}_{B}\cap H|+1=|{\mathbb{X}}_{L_{1}}|+|{\mathbb{X}}_{L_{2}}|+1\geq 2r+3.$
      Thus, we obtain $6r+5\leq|{\mathbb{X}}_{A}|+|{\mathbb{X}}_{B}|=|{\mathbb{X}}|\leq 5r+1$, which is impossible.
    2. (4.2)
      $H\cap K\neq\emptyset$: Because the set ${\mathbb{X}}_{B}$ is not contained in a $4$-plane, the 3-plane $H$ intersects the 2-plane $K$ at a point, say $p$. By Proposition 2.8(c), ${\mathbb{X}}_{K}$ or ${\mathbb{X}}_{K}\cup\{p\}$ has $\mathop{\rm CBP}\nolimits(r-1)$. If ${\mathbb{X}}_{K}$ has $\mathop{\rm CBP}\nolimits(r-1)$ then by Proposition 3.1(a) we have
      $|{\mathbb{X}}_{K}|\geq 2(r-1)+2=2r.$
      When ${\mathbb{X}}_{K}\cup\{p\}$ has $\mathop{\rm CBP}\nolimits(r-1)$, we have
      $|{\mathbb{X}}_{K}|\geq 2(r-1)+1=2r-1.$
      This yields $|{\mathbb{X}}_{K}|\geq 2r-1$. On the other hand, we have ${\mathbb{X}}_{L_{1}}\cup{\mathbb{X}}_{L_{2}}={\mathbb{X}}_{B}\setminus K$, and so this union has $\mathop{\rm CBP}\nolimits(r-2)$. Especially, ${\mathbb{X}}_{L_{1}}$ and ${\mathbb{X}}_{L_{2}}$ also have $\mathop{\rm CBP}\nolimits(r-2)$. By Corollary 2.4, we find $|{\mathbb{X}}_{L_{i}}|\geq(r-2)+2=r$ for $i=1,2$. So, we get
      $|{\mathbb{X}}_{H}|=|{\mathbb{X}}_{L_{1}}|+|{\mathbb{X}}_{L_{2}}|+1\geq 2r+1.$
      This implies $|{\mathbb{X}}_{A}|\geq|{\mathbb{X}}_{H}|+1\geq 2r+2$, and consequently
      $6r+1=(2r+2)+(4r-1)\leq|{\mathbb{X}}_{A}|+|{\mathbb{X}}_{B}|=|{\mathbb{X}}|\leq 5r+1,$
      which is impossible.
  4. Case 5:
    ${\mathbb{X}}_{B}$ is contained in the union of $s$ lines $L_{1},\dots,L_{s}$, where $s=3,4$.
    When $s=3$, three lines $L_{1},L_{2},L_{3}$ must be split, because ${\mathbb{X}}_{B}$ is not contained in a 4-plane. Then $H:=\mathop{\rm span}\nolimits(L_{1},L_{2})$ is a 3-plane, and $H$ and $L_{3}$ are split. In this situation, we go back to Case 2. Now suppose that $s=4$. If any two lines of these lines is contained in a 2-plane then we turn back to either Case 3 or Case 4. This enables us to assume that those lines are pairwise skew. In particular, ${\mathbb{X}}_{B}$ does not lie on a plane configuration of dimension 3. For $i\in\{1,\dots,4\}$, put ${\mathbb{X}}_{L_{i}}:={\mathbb{X}}_{B}\cap L_{i}$ and $H_{ij}:=\mathop{\rm span}\nolimits(L_{i},L_{j})$ for $i\neq j$. Note that ${\mathbb{X}}_{L_{i}}\neq\emptyset$ for $i=1,\dots,4$. Consider the following subcases.
    1. (5.1)
      There are three of four lines that are split. Without loss of generality, we may assume that $L_{1},L_{2}$ and $L_{3}$ are split. Then ${\mathbb{X}}_{L_{1}}\cup{\mathbb{X}}_{L_{2}}\cup{\mathbb{X}}_{L_{3}}={\mathbb{X}}_{B}\setminus L_{4}$ and this union has $\mathop{\rm CBP}\nolimits(r-2)$. By Proposition 2.8(a), the sets of points ${\mathbb{X}}_{L_{1}},{\mathbb{X}}_{L_{2}},{\mathbb{X}}_{L_{3}}$ have $\mathop{\rm CBP}\nolimits(r-2)$. It follows that $|{\mathbb{X}}_{L_{i}}|\geq r$ for $i=1,2,3$ by Corollary 2.4. So, we have
      $|{\mathbb{X}}_{B}|=|{\mathbb{X}}_{L_{1}}|+|{\mathbb{X}}_{L_{2}}|+|{\mathbb{X}}_{L_{3}}|+|{\mathbb{X}}_{L_{4}}|\geq 3r+1.$
      Moreover, since $H_{12}$ is a 3-plane containing $L_{1}$ and $L_{2}$, we also have
      $|{\mathbb{X}}_{A}|\geq|{\mathbb{X}}_{L_{1}}|+|{\mathbb{X}}_{L_{2}}|+1\geq 2r+1.$
      Consequently, in this subcase, we get
      $|{\mathbb{X}}|=|{\mathbb{X}}_{A}|+|{\mathbb{X}}_{B}|\geq 5r+2,$
      which contradicts the hypothesis $|{\mathbb{X}}|\leq 5r+1$.
    2. (5.2)
      All three of four lines are not split. Without lost of generality, we may assume that three lines $L_{1},L_{2},L_{3}$ satisfy the condition that their union ${\mathbb{X}}_{L_{1}}\cup{\mathbb{X}}_{L_{2}}\cup{\mathbb{X}}_{L_{3}}$ contains at least $\frac{3}{4}$ points of ${\mathbb{X}}_{B}$. Also, we have $L_{3}\cap H_{12}\neq\emptyset$.
      • If $L_{3}\subseteq H_{12}$ then we come back to Case 2.
      • If $L_{3}$ intersects $H_{12}$ at a point, then $\mathop{\rm span}\nolimits(H_{12},L_{3})=\mathop{\rm span}\nolimits(L_{1},L_{2},L_{3})$ is a 4-plane. By the choice of the 4-plane $A$, we have
        $|{\mathbb{X}}_{A}|\geq|{\mathbb{X}}_{L_{1}}|+|{\mathbb{X}}_{L_{2}}|+|{\mathbb{X}}_{L_{3}}|\geq\frac{3}{4}|{\mathbb{X}}_{B}|.$
        This implies
        $5r+1\geq|{\mathbb{X}}|=|{\mathbb{X}}_{A}|+|{\mathbb{X}}_{B}|\geq\frac{7}{4}|{\mathbb{X}}_{B}|.$
        Thus, for $r\geq 4$, we get
        $|{\mathbb{X}}_{B}|\leq\frac{20r+4}{7}\leq 4(r-1)+1.$
        By Theorem 1.3, ${\mathbb{X}}_{B}$ lies on a plane configuration of dimension 3, and this is impossible.
Altogether, the proof of the theorem is complete. ∎
Acknowledgements
This work is supported by the Vietnam Ministry of Education and Training under the grant number B2026-DHH-01.